Three coins are selected from nine coins: 3 dimes, 5 nickels, and 1 quarter. In how many ways can the selection be made so that the value of the coins chosen is at least 25 cents?
I couldn't figure out how to use a combination

Answer by edjones(8007)

nCr=Combination of n things taken r at a time.
1C1*8C2=28 One quarter and any other 2 of the other 8 coins.
3C2*5C1=3*5=15 Two dimes and one nickel.
3C3=1 Three dimes
28+16+1=44
.
Ed